OpenStudy (anonymous):

Some one please! check my answre,ty a) write a inequality that presents the interval b) state whether the intesrval is bouded or unbouded. 5x-12>0 a) 12/5

OpenStudy (anonymous):

are there no absolute values here, just \[5x-12>0\]?

OpenStudy (anonymous):

a) |x|<4 b) \(x > \frac{12}{5}\) Looks unbounded to me.

OpenStudy (anonymous):

you lost me

OpenStudy (anonymous):

For the other inequality \( \frac{12}{5} < x < -\frac{12}{5}\) I don't think it is possible

OpenStudy (anonymous):

Not possible I mean not possible

OpenStudy (anonymous):

The other inequality is not possible

OpenStudy (anonymous):

hold on!

OpenStudy (anonymous):

-12/5<x<12/5

OpenStudy (anonymous):

oh that is bounded and possible

OpenStudy (anonymous):

wait wait

OpenStudy (anonymous):

the way i read, there is one question with two parts

OpenStudy (anonymous):

Satellite come back to me :)

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

the question is Let \[12x-5>0\] and then a)write a inequality that presents the interval and b)state whether the intesrval is bouded or unbounded.

OpenStudy (anonymous):

there are two parts to one question right?

OpenStudy (anonymous):

oh I see Nancy don't follow my procedure.

OpenStudy (anonymous):

I did all crap

OpenStudy (anonymous):

so answer is \[12x-5>0\] \[12x>5\] \[x>\frac{5}{12}\] answer is a)\[(\frac{5}{12},\infty)\] and b) it is unbounded

OpenStudy (anonymous):

ty,satellite and ishana94